OpenStudy (anonymous):

Highway Curves If a circular curve without any banking has a radius of R feet, the speed limit L in miles per hour for the curve is L= 1.5"R. (Source: N. Garber.) (a) Find the speed limit for a curve having a radius of 400 feet. (b) If the radius of a curve doubles, what happens to the speed limit? (c) A curve with a 40-mile-per-hour speed limit is being designed. What should be its radius?

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

(a) 1.5*400=600. (b)if radius doubles then speed limit also doubles (c) 40=1.5*R R=40/1.5=26.26